# Performance Trade-offs of High Order Meshless Approximation on Distributed Memory Systems

### v1

@@ Abstract
Real world applications, such as hydrodynamical modeling [1], often require highly detatailed simulations to accurately model systems of interest. These can demand more processing power or memory resources than a single shared memory computer can provide, necessitating the distribution of computations across multiple systems [2]. To distribute workloads across multiple computers, new synchronization methods are required, and existing algorithms designed for shared memory systems must be adapted to function in this distributed environment.
Many physical systems can be described by partial differential equations [3, 4], but few are simple enough to be solved exactly with closed form solutions. Consequently, numerous numerical approaches have been developed to approximate these solutions. In such methods, space must be discretized either by points or volume elements to facilitate computation. One of the possible numerical approaches are meshless methods whose major advantage is the ability to be able to compute solutions on scattered nodes, whereas other methods may require meshing or regular distributions of points with predefined relations.
They solve problems by computing function derivatives at a point by appropriately weighing function values in its immediate neighbourhood. Reliance on a limited number of close points makes them well-suited for distributed memory systems. A part of the point set that discretizes the space can be allocated to each system and then systems exchange values that are needed by a particular system, but are not unique to it [5].
Higher order approximations require more neighbouring nodes to evaluate. Reference [6] has shown that it is advantageous to use higher order approximations in shared memory systems. However it remains unclear whether the same also holds for distributed memory systems where latency and bandwidth limitations constrain the synchronization speed. The aim of this paper is to investigate how the additional complexity impacts the trade-offs of increasing the approximation order and how the distributed algorithm scales.
@@ II Methods
Finite difference methods compute function derivatives on a structured grid of nodes by summing appropriately weighted contributions of values at locations of neighbouring nodes. Meshless methods offer the same capability, but generalised to the unstructured positioning of nodes.
@@
The weights can be computed using the radial basis function-generated finite difference method (RBF-FD) [8] where we obtain them by solving the system
@@ II-A Domain Decomposition
A natural choice for parallelization of our problem is to assign nodes that sample the original domain to subdomains and assign each of these subsets to a compute node which performs the iteration of each step in parallel. The data to be exchanged between the compute nodes are the values of at nodes that are in the support of nodes belonging to a particular subdomain but not contained within it. Therefore, each compute node is assigned a subdomain set which contains nodes that are unique to the subdomain and those that will require values of to be exchanged.
@@
where is the inherent latency of the network, is the amount of data sent and is the bandwidth. To minimise the communication time, we need to minimise the amount of data communicated between the compute nodes. This can be achieved by dividing the domain into subdomains where the nodes are contiguous in the sense that they are in each other’s support. This leads to a collection of subdomain sets that exchange values with neighbouring subdomain sets at their edge nodes.
@@ II-B Implementation
The part of the algorithm concerning local solution of the PDE is implemented in the C++ open source Medusa library [10]. The addition of the domain decomposition is implemented using Open MPI [11], while the parallelisation within the compute node is enabled by OpenMP [12] with #pragma omp parallel for.
Our distributed memory system is a homogeneous cluster of 18 compute nodes connected to a single switch, forming a star topology. Each compute node has two sockets populated with Intel Xeon E5520 CPUs. These have 4 physical cores with 2 threads per core and a single NUMA node per CPU. Each compute node has of memory available.
The program runs with MPI processes (hereinafter referred to as process(es)) (preferably with one process per compute node), where is the number of subdomains. The additional process is referred to as the root and does not contribute to the main iteration, but is only used to collect relevant data from the child processes.
@@ II-C Test Problem
There are many superior methods that can be used to solve the given problem, but the Jacobi method was chosen for its simplicity, as its role is simply to gauge the accuracy of the spatial approximation. It iteratively improves the solution at each node in the interior of the domain from step to step as
@@
Neumann boundary conditions are enforced at the boundary nodes before starting iteration (8). The values of at relevant boundary nodes are computed so that the condition is satisfied. There is no single way to explicitly enforce this condition as satisfying the condition exactly for all relevant boundary nodes at the same time requires the solution of a global system over . For stability reasons and in an effort to keep the domain decomposition simple, we choose to compute the boundary values by considering only the values from the last step at interior nodes in the support of a given boundary node. This method requires updated values at the boundary nodes that are not unique to a subdomain, but are in the support of nodes within the subdomain. Therefore, the original idea of exchanging the required values at nodes from neighbouring subdomains has to be modified to also communicate values at nodes that are in the support of these exterior boundary nodes111This can be seen in Fig. 1 as a perturbation of the exchanged nodes at a subdomain node set’s border..
@@
II-D Benchmarks
Each benchmark run measures the execution time of certain parts of the code. We measure the computation and communication time separately for each process and each step and store the average values for the entire run for each process separately. Since our algorithm does not perform load balancing, care must be taken to synchronise all processes before measuring the communication time, otherwise systematic deviations would occur.
@@ II-D Benchmarks
III Results
The overall performance of the algorithm is shown in Fig. 2, which shows the time per step as a function of the number of nodes discretising the domain for a 16-process configuration. The figure compares how the step time varies depending on the choice of the number of nodes discretising the domain and the choice of the order of augmentation. For domains with or more nodes, the scaling regime is nominal, i.e. increasing the number of nodes corresponds to an increase in the total step time. At lower numbers of nodes, the step times start to approach the order of magnitude of the latency . It was estimated by fitting Eq. 4 to benchmark data of the communication time of individual processes and the amount of data they send to other processes, and was determined to be about . This is only a rough estimate as the communication pattern is likely to be more complex. Firstly, we would need to account for the additional latency introduced by our synchronisation method, as described in Sec. II-D. Secondly, the number of receiver processes varies for each process, so the coordination of communication might introduce additional dynamics that make it difficult to fully account for the delay. Nonetheless, the horizontal line at least serves to indicate the order of communication delay, which could be interpreted as latency.
III-A Strong Scaling
@@
III-B Time and Error Optimization
@@
The right panel of Fig. 4 shows this derived quantity in place of and gives a much clearer performance comparison. It seems like there exist problem sizes where using lower augmentation orders would be advantageous, but such problem sizes would be absurdly small. There also seems that an intersection would be reached between the fourth and sixth order augmentation at low error, but this is just due to the limits of numerical precision. In [6] there is some overlap in performance for different augmentation orders for small problems, but this might be more difficult to achieve on distributed systems due to network latency and the positive effects of increasing the support size discussed later in Sec. III-C. In this case, latency represents a significant amount of time when no processing is taking place, so the advantage of lower order augmentations on small problems which take little time to execute is overshadowed by the time it takes for the network to respond.
III-C Efficiency and Support Size
@@ IV Conclusions
We found that, from accuracy versus computational time perspective, increasing the augmentation order is beneficial. Compared to shared memory computers, where there is some overlap in performance between different approximation orders for small problem sizes, the latency present in communication further suppresses this advantage. This conclusion is further supported by analysis of compute and communication time scaling with support size where the latter scales slower than the former, ensuring that efficiency increases when more accurate solutions are required, as long as the radius of the support remains below the order of the size of the subdomain.
@@ acknowledgment
The authors would like to acknowledge the financial support of the Slovenian Research and Innovation Agency (ARIS) research core funding No. P2-0095 and the Young Researcher programmes PR-10468 and PR-13389.

### v2

@@ Abstract
Real world applications, such as hydrodynamical modeling [1], often require highly detatailed simulations to accurately model systems of interest. These can demand more processing power or memory resources than a single shared memory computer can provide, necessitating the distribution of computations across multiple systems [2]. To distribute workloads across multiple computers, new synchronization methods are required, and existing algorithms designed for shared memory systems must be adapted to function in parallel due to the distributed nature of the high-performance computing (HPC) systems.
Many physical systems can be described by partial differential equations, but few are simple enough to be solved exactly with closed form solutions. Consequently, numerous numerical approaches have been developed to approximate these solutions. In such methods, space must be discretized either by points or volume elements to facilitate computation. One of the possible numerical approaches are meshless methods whose major advantage is the ability to be able to compute solutions on scattered nodes, whereas other methods may require meshing or regular distributions of points with predefined relations.
Meshless methods approximate function derivatives at a point by appropriately weighing function values in its immediate neighbourhood. Reliance on a limited number of close points makes them well-suited for distributed memory systems. A part of the point set that discretizes the space can be allocated to each system and then systems exchange values that are needed by a particular system, but are not unique to it [3].
Higher order approximations provide better rate of convergence111With increasing discretization accuracy, i.e. decreasing distance between nodes, a higher order approximation will approach the exact value faster. but require more neighbouring nodes to evaluate. For the radial basis function-generated finite differences (RBF-FD) method [4] it has already been shown by Jančič et al. [5] that the increased rate of convergence of the higher order approximation outweighs the additional computational costs when high accuracy is required. However, it remains unclear whether the same also holds for distributed memory systems where latency and bandwidth limitations impose additional communication cost. The aim of this paper is to investigate how the additional complexity impacts the trade-offs of increasing the approximation order and how the distributed algorithm scales.
@@ II Methods
Finite difference methods [6] compute function derivatives on a structured grid of nodes by summing appropriately weighted contributions of values at locations of neighbouring nodes. Meshless methods offer the same capability, but generalised to the unstructured positioning of nodes.
@@
The weights can be computed using the RBF-FD [4] method where we obtain them by solving the system
@@ II-A Domain Decomposition
A natural choice for parallelization of our problem is to assign nodes (as described at the beginning of Sec. II) that sample the original domain to subdomains and assign each of these subsets to a compute node (CN) which performs the iteration of each step in parallel. The data to be exchanged between the CNs are the values of at nodes that are in the support of nodes belonging to a particular subdomain but not contained within it. Therefore, each CN is assigned a subdomain set which contains nodes that are unique to the subdomain and those that will require values of to be exchanged.
@@
where is the inherent latency of the network, is the amount of data sent and is the bandwidth. Since and are properties of the underlying hardware and thus out of our control, the only way for us to affect the communication time is by reducing the amount of data communicated between the CNs. Communication minimisation can be achieved by dividing the domain into subdomains where the nodes are contiguous in the sense that they are in each other’s support. This leads to a collection of subdomain sets that exchange values with neighbouring subdomain sets at their edge nodes.
@@ II-B Implementation
The part of the algorithm concerning local solution of the PDE is implemented in the C++ open source Medusa library [9]. The addition of the domain decomposition is implemented using Open MPI [10], while the parallelisation within the CN is enabled by OpenMP [11] with #pragma omp parallel for.
Our distributed memory system is a homogeneous cluster of 18 CNs connected with gigabit ethernet to a single switch, forming a star topology. Each CN has two sockets populated with Intel Xeon E5520 CPUs. These have 4 physical cores with 2 threads per core and a single NUMA node per CPU. Each CN has of memory available.
The program runs with MPI processes (hereinafter referred to as process(es)) (preferably with one process per CN), where is the number of subdomains. The additional process is referred to as the root and does not contribute to the main iteration, but is only used to collect relevant data from the child processes.
@@ II-C Test Problem
There are many superior methods that can be used to solve the given problem, but the Jacobi method [12] was chosen for its simplicity, as its role is simply to gauge the accuracy of the underlying spatial derivative approximation. Edge node communication times are proportional for any method using RBF-FD approximation of the same order of accuracy, making these results generalizable. The method iteratively improves the solution at each node in the interior of the domain from step to step as
@@
Neumann boundary conditions are enforced at the boundary nodes before starting iteration (8). The values of at relevant boundary nodes are computed so that the condition is satisfied. There is no single way to explicitly enforce this condition as satisfying the condition exactly for all relevant boundary nodes at the same time requires the solution of a global system over . For stability reasons and in an effort to keep the domain decomposition simple, we choose to compute the boundary values by considering only the values from the last step at interior nodes in the support of a given boundary node. This method requires updated values at the boundary nodes that are not unique to a subdomain, but are in the support of nodes within the subdomain. Therefore, the original idea of exchanging the required values at nodes from neighbouring subdomains has to be modified to also communicate values at nodes that are in the support of these exterior boundary nodes222This can be seen in Fig. 1 as a perturbation of the exchanged nodes at a subdomain node set’s border..
@@
III Results
III-A Timing setup
Each benchmark run of the solution to the Poisson problem described in Sec. II-C measures the execution time of certain parts of the code. We measure the computation and communication time separately for each process and each step and store the average values for the entire run for each process separately. Since our algorithm does not perform load balancing, care must be taken to synchronise all processes before measuring the communication time, otherwise systematic deviations would occur.
@@ III-A Timing setup
III-B First Results
The overall performance of the algorithm is shown in Fig. 2, which shows the time per step as a function of the number of nodes discretising the domain for a 16-process configuration. The figure compares how the step time varies depending on the choice of the number of nodes discretising the domain and the choice of the order of augmentation. For domains with or more nodes, the scaling regime is nominal, i.e. increasing the number of nodes corresponds to an increase in the total step time. At lower numbers of nodes, the step times start to approach the order of magnitude of the latency . It was estimated by fitting Eq. 4 to benchmark data of the communication time of individual processes and the amount of data they send to other processes, and was determined to be about . This is only a rough estimate as the communication pattern is likely to be more complex. Firstly, we would need to account for the additional latency introduced by our synchronisation method, as described in Sec. III-A. Secondly, the number of receiver processes varies for each process, so the coordination of communication might introduce additional dynamics that make it difficult to fully account for the delay. Nonetheless, the horizontal line at least serves to indicate the order of communication delay, which could be interpreted as latency.
III-C Strong Scaling
@@
III-D Time and Error Optimization
@@
The right panel of Fig. 4 shows this derived quantity in place of and gives a much clearer performance comparison. It seems like there exist problem sizes where using lower augmentation orders would be advantageous, but such problem sizes would be absurdly small. There also seems that an intersection would be reached between the fourth and sixth order augmentation at low error, but this is just due to the limits of numerical precision. In [5] there is some overlap in performance for different augmentation orders for small problems, but this might be more difficult to achieve on distributed systems due to network latency and the positive effects of increasing the support size discussed later in Sec. III-E. In this case, latency represents a significant amount of time when no processing is taking place, so the advantage of lower order augmentations on small problems which take little time to execute is overshadowed by the time it takes for the network to respond.
III-E Efficiency and Support Size
@@ IV Conclusions
We found that, from accuracy versus computational time perspective, increasing the augmentation order is beneficial. Compared to shared memory computers, where there is some overlap in performance between different approximation orders for small problem sizes, the latency present in communication further suppresses this advantage. This conclusion is further supported by analysis of compute and communication time scaling with support size where the latter scales slower than the former, ensuring that efficiency increases when more accurate solutions are required, as long as the radius of the support remains below the order of the size of the subdomain. Therefore we conclude that on distributed memory computers it is beneficial to use higher order methods, regardless of the problem size, as long as the error introduced by finite difference arithmetic [5] isn’t significant.
@@ acknowledgment
The authors would like to acknowledge the financial support of the Slovenian Research Agency (ARRS) research core funding No. P2-0095 and the Young Researcher programmes PR-10468 and PR-13389.